D = 50 - 0.5P and S= 20 + 0.25P

Calculate Quantity Demanded when price is 10
Calculate Quantity Demanded when price is 20
Calculate the equilibrium prices and quantities
Calculate the Shortage / Surplus if government imposes a regulatory price of Rs. 60
1
Expert's answer
2020-03-25T10:54:00-0400

"D=50\u22120.5P" ,"S=20+0.25P"

a. Qd when P = $10,

"Qd = 50 - 0.5 \times10" = 45 units.

b. Qs when P = $20

"Qs = 20 + 0.25 \times 20" = 25 units.

c. At equilibrium, Qd = Qs,

"Qe=45+\frac{25}{2}" =35 units

P= $15

d. shortage /surplus

"D=50 -0.5\times60" =20 units

"S=20+0.25\times60" =35 units